OpenStudy (woodrow73):

Would having 2 220 ohm resistors in series be equivalent to the affect of a single 440 ohm resistor?

OpenStudy (anonymous):

Yes. Cuase the equivalent resistance, in a series combination is Req = R1 + R2 ...

OpenStudy (anonymous):

The resistance is equivalent

OpenStudy (anonymous):

Hence, here, Req = 220 + 220 = 440 ohm which will work

OpenStudy (woodrow73):

Thanks! And if you feed a current/voltage into the cathode of a resistor, am I to believe that no current would ever travel through the anode? (cathode = -) (anode = +)

OpenStudy (anonymous):

But if you consider from the perspective of each resistor, they're not necessarily going to get the same treatment as the 440 resistor got.

OpenStudy (anonymous):

They have to share the volts

OpenStudy (woodrow73):

I still haven't learned current vs. voltage.. I'll study up on them now.

OpenStudy (anonymous):

Like if you had 5V, the 440 ohm will experience all of it, but the 220 ohm will each experience 2.5V

OpenStudy (woodrow73):

Just ohm's law - and using that to calculate acceptable current for led's and stuff.

OpenStudy (anonymous):

I bring this up because if the resistors are expecting a certain voltage range, then there are consequences

OpenStudy (woodrow73):

hmm- interesting. I'll definitely keep that in mind.

OpenStudy (anonymous):

I wasn't sure if you were doing a homework question or a legitimate project, that's why I bring it up

OpenStudy (woodrow73):

Are you at all familiar with arduino-ish technology's? And no, just trying to tackle these concepts both for my understanding & so I can better handle the Arduino.

OpenStudy (anonymous):

Not really

OpenStudy (woodrow73):

Thanks a bunch for all the help @wio & @Abhilash11